OpenStudy (anonymous):

Transform this equation to slope-intercept form: 6x + 4y = 8

OpenStudy (hysenberg):

y=mx+c write in this form

OpenStudy (kkutie7):

Y=2-3/2x

OpenStudy (hysenberg):

6x + 4y = 8 y=8/4-6/8*x y=2-1.5x

OpenStudy (kkutie7):

^^^6/8???? It should be 6/4

OpenStudy (hysenberg):

yeAH